# Operation of the Trigger System for the ICARUS Detector at Fermilab

### v1

@@ 1 Introduction
The detector consists of two identical modules, $19$m long, $3.6$m wide and $3.9$m tall each, filled with 760 t of ultra-pure liquid argon. A detailed detector description can be found in [2]. Each module houses two LAr TPCs separated by a 58%-transparent shared cathode with a maximum drift distance of $1.5$m, equivalent to $\approx\!1$ms drift time for the nominal 500 V/cm electric drift field. The TPCs (53,248 wires in total) are characterized by three parallel readout wire planes with $3$mm pitch, placed $3$mm apart from each other and oriented at 0 and ±60° with respect to the horizontal direction. By appropriate voltage biasing, the first two planes (Induction 1 and Induction 2) provide a nondestructive charge measurement, whereas the ionization charge is fully collected by the last (Collection) plane.
@@
The main ICARUS trigger exploits the coincidence of the BNB and NuMI beams spills, $1.6$µs and $9.5$µs respectively, with the prompt scintillation light signals detected by the PMT system. Due to the $<2$GeV ($<5$GeV) neutrino energy range of BNB (NuMI) beams (see figure 1), the bulk of neutrino interaction events is expected to be contained in about 4 m along the beam direction, supporting the choice of a trigger based on the identification of fired PMTs inside a limited TPC region [4]. The data taking is complemented by routinely collecting events without any request on the scintillation light, with and without the presence of the beam, and also with high-rate external triggers for calibration purposes.
@@
The present paper is devoted to the description and characterization of the trigger system deployed for the first ICARUS data taking for physics with BNB and NuMI neutrino beam in the periods June 9 – July 14, 2022 (“Run1”) and December 20, 2022 – June 10, 2023 (“Run2”). The instrumentation adopted for the trigger signal generation starting from the PMT signals distribution, the accelerator signals of proton on target extraction and the required time synchronization system are described in section 2, as well as the light-based triggering logic and the specific triggers implemented. A detailed study of the trigger performance in terms of event recognition efficiency is performed with cosmic rays, as described in section 3.
@@ 2 The ICARUS trigger system
The trigger architecture allows for the acquisition of different types of events, with and without the presence of the neutrino beams (see figure 2). The main ICARUS global trigger (“on-beam majority”) is based on the multiplicity of PMT signals in coincidence within a “beam gate” opened at the expected arrival time of neutrinos as provided by the proton spill extraction signals of BNB and NuMI beams (see section 2.2). Off-beam cosmic ray events are collected with similar “off-beam majority” global triggers, based on the coincidence between light signals and off-beam gates generated 33 ms after each on-beam gate, away from the beam spills. In addition, on-beam and off-beam “minimum bias” global triggers are generated in the presence of the corresponding gates, regardless of the scintillation light signals, to provide an unbiased data sample. Minimum-bias triggers are collected once every 20 off-beam gates, and every 200 (60) on-beam gates for BNB (NuMI), to minimize the impact on triggered beam data taking, but to still collect an adequate sample of unbiased triggers. Unbiased data can be used to validate the performance of the majority trigger, for detector physics studies, and calibrations. The ICARUS trigger system can also operate in a multi-Hz “calibration” mode when the accelerators are not operational, opening gates synchronized to a pulse generator, to collect off-beam data at a much higher rate than during beam operations.
@@ 2.1 Hardware setup and implementation
The scintillation light is collected by 360 Hamamatsu R5912-MOD 8” PMTs made sensitive to vacuum ultraviolet (VUV) photons by 200 µg/cm2 tetraphenyl butadiene coating deposited on their external window [4]. The PMTs are installed behind the wire planes of each of the 4 TPCs, with 90 PMTs per TPC (figure 3).
@@
The signals coming from the 360 PMTs are read out by 24 CAEN V1730B digitizers, installed in 8 VME crates, 3 boards per crate. Each digitizer hosts 16 flash ADC channels, with 14-bit 500-MS/s sampling and 2 Vpp input dynamic range. In each board 15 channels are used for the acquisition of PMT signals, while the 16th channel is used to record three reference timing signals: the global trigger signal, the proton spill extraction signal, also used to generate beam gates (section 2.2), and the signal of actual arrival of protons on the beam target from the resistive wall monitor counters (RWM) [6].
@@
The digitizers generate trigger-request logical patterns via low voltage differential signal outputs (LVDS) which are activated by the presence of PMT signals with amplitude exceeding a digitally programmed threshold. Each digitizer produces 8 LVDS signals, 7 OR combination of PMT pairs and one PMT singlet, that are fed into the hardware modules that implement the trigger logic.
@@
Figure 4: Schematics of the trigger system architecture layout and the main SPEXI functions as indicated in the circle. (right) Photo of the NI-1082 PXIe crate installed in a rack on the east mezzanine of the SBN-FD building.
@@
The trigger hardware setup (figure 4) consists of a single NI PXIe-1082 crate containing a NI PXIe-8840 real time controller (RTC), one mezzanine carrier board in PXIe form factor by INCAA Computers (SPEXI) and three NI PXIe-7820R Field-Programmable Gate Arrays (FPGAs). The SPEXI board generates signals based on the information of the proton extraction at BNB or NuMI through the White Rabbit (WR) network (section 2.2). Such signals are: (i) a pulse-per-second (PPS) used both to reset the PMT digitizer time counters and as one of the reference signals of the CRT Front-End Boards [3]; (ii) $62.5$MHz clock for the PMT digitizers; (iii) $10$MHz clock and trigger for the TPC digitizers via a serial link named TTLink [7]; (iv) $2.2$µs-wide ($10.1$µs-wide) beam gate for BNB (NuMI); (v) $2$ms-wide enable gate for activating the PMT waveform readout outside of the beam spill. While the clock signals are output on the front panel, the beam gate signals are distributed through the back-plane of the trigger crate to be available to the other modules for further use.
@@
Two of the three FPGAs, one per cryostat, are dedicated to processing the LVDS signals from the PMT digitizer boards to generate PMT trigger primitives according to the logic described in section 2.3. Each trigger primitive is distributed both on the back-plane of the trigger crate and to the PMT digitizer boards for PMT signal recording when in coincidence with the beam enable gate window. The third FPGA combines all signals present on the back-plane (beam gates and PMT trigger primitives) to produce the global trigger, and vetoes the generation of other trigger primitives for the rest of the beam gate to ensure having only one global trigger for each spill.
@@
When a PMT digitizer receives a trigger primitive, it generates another timestamp (trigger time tag, TTT) that is written in the header of the PMT data fragments. The TTT comes from an internal 31-bit counter, resets every second to avoid uncontrolled overflows using the PPS signal from the SPEXI, and is characterized by a $16$ns time resolution. As a result, the timestamp associated to the PMT waveforms represents the number of nanoseconds elapsed since the last full second of GPS time. In order to improve the synchronization between trigger and PMT readout, a copy of the global trigger signal is sampled in ${}^{\text{th}}$ channel of one digitizer in each crate, and its time is used as an external reference to correct the timestamps event by event. The time of the PMT pulses is then reconstructed by combining together the timestamp of the global trigger from the White Rabbit system and the TTT, corrected by the relative time of the global trigger signal within the digitized window.
@@ 2.2 Accessing the proton accelerator signals
Distribution from the beam locations to the ICARUS building (SBN-FD) occurs via White Rabbit (WR), an Ethernet based network for synchronization of distributed systems with sub-nanosecond accuracy and precision better than $50$ps [9]. It consists of one master switch, disciplined by a GPS-locked atomic clock, connected to three nodes: 2 PCI-e cards (SPEC) in computers at locations MI-12 and MI-60, and the aforementioned SPEXI card in the NI crate at SBN-FD (see figure 5).
@@
The SPEC WR nodes perform the timestamping of the leading edge of all the early warning signals and broadcast them over the WR network with $8$ns resolution using Digital Input Output (DIO) mezzanine and Network Interface Card (NIC) core. The SPEXI receives such information and uses it, with appropriate time offsets, to open the enable and beam gates. These time offsets were validated and fine-tuned with beam data. The duration of the beam gates (2.2 and $10.1$µs for BNB and NuMI) is programmed to be larger than the duration of the beam extraction (1.6 and $9.5$µs respectively) to account for possible jitters. The excess of neutrinos and beam-associated particles over the cosmic rays, clearly identified inside the trigger beam gates for both BNB and NuMI in the collected data (figure 6), confirms the synchronization of the trigger gates with the neutrino beam spills.
@@ 2.3 Implementation of the PMT-majority logic
The longitudinal span of each TPC was divided into three logical windows, each one $6$m long and containing the signals of 60 PMTs (30 PMTs per TPC, on opposite sides of the central cathode), in which the multiplicity of OR-paired PMT signals exceeding a 390 ADC threshold (approximately 13 photoelectrons) is evaluated. This initial configuration, deployed in Run1, was upgraded in Run2 by implementing two additional overlapping windows, shifted by half a window (figure 7), to improve the trigger uniformity along the beam direction.
Figure 7: (left) View from the top of an ICARUS module, with two LAr-TPCs separated by a common cathode. For the trigger logic, light signals are evaluated in three side-by-side $6$m longitudinal module slices. For Run2, two additional slices (dashed lines) were introduced to improve efficiency and uniformity. (right) Side-view of the module. In the highlighted window, containing 60 PMTs (30 per TPC), an example of a Mj=5 pattern is shown, with PMTs paired into LVDS channels.
@@
For each $6$m window, the multiplicity of LVDS channels from PMT pairs is evaluated on every $25$ns FPGA clock cycle, independently by the two FPGAs dedicated to the east and west modules. When at least 5 LVDS channels are fired within one window and inside the beam gate, a local PMT trigger primitive is generated in the corresponding module (figure 8) to evaluate the global trigger decision. The choice of the majority levels was driven by the need of a high trigger efficiency for collecting BNB and NuMI neutrinos, while ensuring a sustainable data rate and size. In particular, looser conditions (at least 5 active LVDS channels, “Mj=5”) were applied inside the beam gate to maximize the sensitivity to neutrino interactions with energy depositions as low as 200 MeV. Tighter conditions, at least 10 and 9 active LVDS channels (Mj=10 and Mj=9), were applied outside of the spills in Run1 and Run2 respectively, to limit the amount of collected PMT data while still being able to reject energetic cosmic rays in the analysis stage.
For the Run2 data taking, a typical trigger rate of $\approx 0.7$Hz was obtained, including $\approx 0.2$Hz and $\approx 0.25$Hz respectively from the BNB and NuMI majority components and $\approx 0.25$Hz from minimum-bias triggers (table 1). This resulted in a manageable data readout bandwidth with good operational stability.
@@ 2.4 Collection of the triggered events
Among other information, the trigger packet contains the gate opening and trigger timestamps, the gate type (on-beam or off-beam), the trigger type (majority or minimum bias trigger), and the state of each of the LVDS channels at trigger time, providing PMT-level information on where the trigger occurred within the detector. The inclusion of the $8$ns-precise timestamps from the White Rabbit system allows to properly align the beam gate with the beam spills (see figure 6).
@@ 3.1 Timing of cosmic ray tracks in the TPC
The spatial and time correspondence of CRT signals with PMT light and TPC charge signals are used to reconstruct the time and position of the cosmic rays based on the corresponding signal on the CRT (“hit”). For this purpose, the trajectory of each track is linearly extrapolated using its direction at both track ends to intersect the CRT counters. Among the CRT hits whose time is compatible with the track position in the TPC, the hit with the smallest distance on the CRT plane ($d_{\text{CRT}}$) to any one of the two track-CRT intersections is chosen and its time is assigned to the track. When applying the matching to the collected data, 76% of the cosmic ray tracks are matched to the Top-CRT, and 24% to the Side CRT [^8] The matching with $d_{\text{CRT}}<1\,\text{m}$ has an estimated efficiency of $\approx 90\%$ for the Top CRT and $\approx 50\%$ for the Side CRT [15].(figure 9, right).
The PMT signal interval used to evaluate the response of the trigger to the track is determined by the time of the matched CRT hit. The distribution of recorded times for CRT hits and PMT signals with respect to the digitized global trigger signal is shown in figure 9 (left).
@@ 3.2 Trigger emulation
The trigger response was studied with a software emulation of the trigger hardware operations based on the recorded PMT waveforms, to evaluate different trigger conditions with a single sample of unbiased data. The emulation algorithm was applied in a $300$ns window around the time of reconstructed tracks (section 3.1) to include the associated prompt scintillation light signals. The algorithm replicated trigger hardware operations: discrimination of PMT signals at the leading edge and their pairing in OR logic, clustering in 3 or 5 logic windows (figure 7), counting of PMT-pairs to meet the required majority setting, and inferring the time the trigger would fire at.
The consistency between emulation and hardware operation was validated on an event-by-event basis with a dedicated sample of minimum bias data that also included the information of the PMT-majority trigger at the hardware level. The emulation was found to miss a trigger in less than 0.1% of the hardware-triggered events, due to the unavoidable presence of small signal jitters in the hardware. Conversely, the hardware trigger confirmed the emulated triggers at the level of $0.6\%$ (figure 10) after a $\approx 5\%$ correction was introduced to fix an inconsistency between the misconfigured hardware and the software emulation, which discriminated PMT signals on the trailing and leading edge respectively (see section 3.4).
Figure 10: Emulated triggers in $<20\ \text{h}$ minimum bias runs, after the correction to the PMT readout is introduced, so that PMT signals are discriminated at the leading edge in both the emulation and the hardware. The fiducialized $[0.2,1.9]$µs beam gate window is highlighted.

### v2

@@ 1 Introduction
The detector consists of two identical modules, $19$ m long, $3.6$ m wide and $3.9$ m tall each, filled with 760 t of ultra-pure liquid argon. A detailed detector description can be found in [2]. Each module houses two LAr TPCs separated by a 58%-transparent shared cathode with a maximum drift distance of $1.5$ m, equivalent to $\approx\!1$ ms drift time for the nominal 500 V/cm electric drift field. The four TPCs (53,248 wires in total) are each characterized by three parallel readout wire planes with $3$ mm pitch, placed $3$ mm apart from each other and oriented at 0 and ±60° with respect to the horizontal direction. By appropriate voltage biasing, the first two planes (Induction 1 and Induction 2) provide a nondestructive charge measurement, whereas the ionization charge is fully collected by the last (Collection) plane.
@@
The main ICARUS trigger exploits the coincidence of the BNB and NuMI beams spills, $1.6$ µs and $9.5$ µs respectively, with the prompt scintillation light signals detected by the PMT system. Due to the <2<2 GeV (<5<5 GeV) neutrino energy range of BNB (NuMI) beams (see figure˜1), the bulk of neutrino interaction events is expected to be contained in about 4 m along the beam direction, supporting the choice of a trigger based on the identification of fired PMTs inside a limited TPC region [4]. The data taking is complemented by routinely collecting events without any request on the scintillation light, with and without the presence of the beam, and also with high-rate external triggers for calibration purposes.
@@
The present paper is devoted to the description and characterization of the trigger system deployed for the first ICARUS data taking for physics with BNB and NuMI neutrino beam in the periods June 9 – July 14, 2022 (“Run1”) and December 20, 2022 – June 10, 2023 (“Run2”). The instrumentation adopted for the trigger signal generation starting from the PMT signals distribution, the accelerator signals of proton on target extraction and the required time synchronization system are described in section˜2, as well as the light-based triggering logic and the specific triggers implemented. A detailed study of the trigger performance in terms of event recognition efficiency is performed with cosmic rays, as described in section˜3.
@@ 2 The ICARUS trigger system
The trigger architecture allows for the acquisition of different types of events, with and without the presence of the neutrino beams (see figure˜2). The main ICARUS global trigger (“on-beam majority”) is based on the multiplicity of PMT signals in coincidence within a “beam gate” opened at the expected arrival time of neutrinos as provided by the proton spill extraction signals of BNB and NuMI beams (see section˜2.2). Off-beam cosmic ray events are collected with similar “off-beam majority” global triggers, based on the coincidence between light signals and off-beam gates generated 33 ms after each on-beam gate, away from the beam spills. In addition, on-beam and off-beam “minimum bias” global triggers are generated in the presence of the corresponding gates, regardless of the scintillation light signals, to provide an unbiased data sample. Minimum-bias triggers are collected once every 20 off-beam gates, and every 200 (60) on-beam gates for BNB (NuMI), to minimize the impact on triggered beam data taking, but to still collect an adequate sample of unbiased triggers. Unbiased data can be used to validate the performance of the majority trigger, for detector physics studies, and calibrations. The ICARUS trigger system can also operate in a multi-Hz “calibration” mode when the accelerators are not operational, opening gates synchronized to a pulse generator, to collect off-beam data at a much higher rate than during beam operations.
@@ 2.1 Hardware setup and implementation
The scintillation light is collected by 360 Hamamatsu R5912-MOD 8” PMTs made sensitive to vacuum ultraviolet (VUV) photons by 200 µg/cm2 tetraphenyl butadiene coating deposited on their external window [4]. The PMTs are installed behind the wire planes of each of the 4 TPCs, with 90 PMTs per TPC (figure˜3).
@@
The signals coming from the 360 PMTs are read out by 24 CAEN V1730B digitizers, installed in 8 VME crates, 3 boards per crate. Each digitizer hosts 16 flash ADC channels, with 14-bit 500-MS/s sampling and 2 V input dynamic range. In each board 15 channels are used for the acquisition of PMT signals, while the 16th channel is used to record three reference timing signals: the global trigger signal, the proton spill extraction signal, also used to generate beam gates (section˜2.2), and the signal of actual arrival of protons on the beam target from the resistive wall monitor counters [6]. This last signal enables physics analyses to determine the crossing time of beam neutrino bunches with nanosecond precision [7].
@@
The digitizers generate trigger-request logical patterns via low voltage differential signal outputs (LVDS), which are activated by the presence of PMT signals with amplitude exceeding a digitally programmed threshold. Each digitizer produces 8 LVDS signals, 7 OR combinations of PMT pairs and one PMT singlet, that are fed into the hardware modules that implement the trigger logic.
@@
Figure 4: (Left) Schematics of the trigger system architecture layout and the main SPEXI functions as indicated in the expanded view. (Right) Photo of the NI-1082 PXIe crate installed in a rack on the east mezzanine of the SBN-FD building.
@@
The trigger hardware setup (figure˜4) consists of a single NI PXIe-1082 crate containing a NI PXIe-8840 real time controller (RTC), one mezzanine carrier board in PXIe form factor by INCAA Computers (SPEXI) and three NI PXIe-7820R Field-Programmable Gate Arrays (FPGAs). The SPEXI board generates signals based on the information of the proton extraction at BNB or NuMI through the White Rabbit (WR) network (section˜2.2). Such signals are: (i) a pulse-per-second (PPS) used both to reset the PMT digitizer time counters and as one of the reference signals of the CRT Front-End Boards [3]; (ii) $62.5$ MHz clock for the PMT digitizers; (iii) $10$ MHz clock and trigger for the TPC digitizers via a serial link named TTLink [8]; (iv) $2.2$ µs-wide ($10.1$ µs-wide) beam gate for BNB (NuMI); (v) $2$ ms-wide enable gate for activating the PMT waveform readout outside of the beam spill. While the clock signals are output on the front panel, the beam gate signals are distributed through the back-plane of the trigger crate to be available to the other modules for further use.
@@
Two of the three FPGAs, one per cryostat, are dedicated to processing the LVDS signals from the PMT digitizer boards to generate PMT trigger primitives according to the logic described in section˜2.3. Each trigger primitive is distributed both on the back-plane of the trigger crate and to the PMT digitizer boards for PMT signal recording when in coincidence with the beam enable gate window. The third FPGA combines all signals present on the back-plane (beam gates and PMT trigger primitives) to produce the global trigger, and vetoes the generation of other trigger primitives for the rest of the beam gate to ensure having only one global trigger for each spill.
@@
[…]
@@ 2.2 Accessing the proton accelerator signals
Distribution from the beam locations to the ICARUS building (SBN-FD) occurs via White Rabbit (WR), an Ethernet based network for synchronization of distributed systems with sub-nanosecond accuracy and precision better than $50$ ps [10]. It consists of one master switch, disciplined by a GPS-locked atomic clock, connected to three nodes: 2 PCI-e cards (SPEC) in computers at locations MI-12 and MI-60, and the aforementioned SPEXI card in the NI crate at SBN-FD (see figure˜5).
@@
The SPEC WR nodes perform the timestamping of the leading edge of all the early warning signals and broadcast them over the WR network with $8$ ns resolution using Digital Input Output mezzanine (FMC-DIO) and Network Interface Card (NIC) core. The SPEXI receives such information and uses it, with appropriate time offsets, to open the enable and beam gates. These time offsets were validated and fine-tuned with beam data. The duration of the beam gates (2.2 and $10.1$ µs for BNB and NuMI) is programmed to be larger than the duration of the beam extraction (1.6 and $9.5$ µs respectively) to account for possible jitters. The excess of neutrinos and beam-associated particles over the cosmic rays, clearly identified inside the trigger beam gates for both BNB and NuMI in the collected data (figure˜6), confirms the synchronization of the trigger gates with the neutrino beam spills.
@@ 2.3 Implementation of the PMT-majority logic
The longitudinal span of each TPC was divided into three logical windows, each one $6$ m long and containing the signals of 60 PMTs (30 PMTs per TPC, on opposite sides of the central cathode), in which the multiplicity of OR-paired PMT signals exceeding a 390 ADC threshold (approximately 13 photoelectrons) is evaluated. This initial configuration, deployed in Run1, was upgraded in Run2 by implementing two additional overlapping windows, shifted by half a window (figure˜7), to improve the trigger uniformity along the beam direction.
Figure 7: (left) View from the top of an ICARUS module, with two LAr-TPCs separated by a common cathode. For the trigger logic, light signals are evaluated in three side-by-side 66 m longitudinal module slices. For Run2, two additional slices (dashed lines) were introduced to improve efficiency and uniformity. (right) Side-view of the module. In the highlighted window, containing 60 PMTs (30 per TPC), an example of a Mj=5 pattern is shown, with PMTs paired into LVDS channels.
@@
For each $6$ m window, the multiplicity of LVDS channels from PMT pairs is evaluated on every $25$ ns FPGA clock cycle, independently by the two FPGAs dedicated to the east and west modules. When at least 5 LVDS channels are fired within one window and inside the beam gate, a local PMT trigger primitive is generated in the corresponding module (figure˜8) to evaluate the global trigger decision. The choice of the majority levels was driven by the need of a high trigger efficiency for collecting BNB and NuMI neutrinos, while ensuring a sustainable data rate and size. In particular, looser conditions (at least 5 active LVDS channels, “Mj=5”) were applied inside the beam gate to maximize the sensitivity to neutrino interactions with energy depositions as low as 200 MeV. Tighter conditions, at least 10 and 9 active LVDS channels (Mj=10 and Mj=9), were applied outside of the spills in Run1 and Run2 respectively, to limit the amount of collected PMT data while still being able to reject energetic cosmic rays in the analysis stage.
For the Run2 data taking, a typical trigger rate of $\approx 0.7$ Hz was obtained, including $\approx 0.2$ Hz and $\approx 0.25$ Hz respectively from the BNB and NuMI majority components and $\approx 0.25$ Hz from minimum-bias triggers (table˜1). This resulted in a manageable data readout bandwidth with good operational stability.
@@ 2.4 Collection of the triggered events
Among other information, the trigger packet contains the gate opening and trigger timestamps, the gate type (on-beam or off-beam), the trigger type (majority or minimum bias trigger), and the state of each of the LVDS channels at trigger time, providing PMT-level information on where the trigger occurred within the detector. The inclusion of the $8$ ns-precise timestamps from the White Rabbit system allows to properly align the beam gate with the beam spills (see figure˜6).
@@ 3.1 Timing of cosmic ray tracks in the TPC
The spatial and time correspondence of CRT signals with PMT light and TPC charge signals are used to reconstruct the time and position of the cosmic rays based on the corresponding signal on the CRT (“hit”). For this purpose, the trajectory of each track is linearly extrapolated using its direction at both track ends to intersect the CRT counters. Among the CRT hits whose time is compatible with the track position in the TPC, the hit with the smallest distance on the CRT plane ($d_{\text{CRT}}$) to any one of the two track-CRT intersections is chosen and its time is assigned to the track. When applying the matching to the collected data, 76% of the cosmic ray tracks are matched to the Top-CRT, and 24% to the Side CRT [^8] The matching with $d_{\text{CRT}}<1\,\text{m}$ has an estimated efficiency of $\approx 90\%$ for the Top CRT and $\approx 50\%$ for the Side CRT [16].(figure˜9, right).
The PMT signal interval used to evaluate the response of the trigger to the track is determined by the time of the matched CRT hit. The distribution of recorded times for CRT hits and PMT signals with respect to the digitized global trigger signal is shown in figure˜9 (left).
@@ 3.2 Trigger emulation
The trigger response was studied with a software emulation of the trigger hardware operations based on the recorded PMT waveforms, to evaluate different trigger conditions with a single sample of unbiased data. The emulation algorithm was applied in a $300$ ns window around the time of reconstructed tracks (section˜3.1) to include the associated prompt scintillation light signals. The algorithm replicated trigger hardware operations: discrimination of PMT signals at the leading edge and their pairing in OR logic, clustering in 3 or 5 logic windows (figure˜7), counting of PMT-pairs to meet the required majority setting, and inferring the time the trigger would fire at.
The consistency between emulation and hardware operation was validated on an event-by-event basis with a dedicated sample of minimum bias data that also included the information of the PMT-majority trigger at the hardware level. The emulation was found to miss a trigger in less than 0.1% of the hardware-triggered events, due to the unavoidable presence of small signal jitters in the hardware. Conversely, the hardware trigger confirmed the emulated triggers at the level of $0.6\%$ (figure˜10) after a $\approx 5\%$ correction was introduced to fix an inconsistency between the misconfigured hardware and the software emulation, which discriminated PMT signals on the trailing and leading edge respectively (see section˜3.4).
Figure 10: Emulated triggers in <20h<20\ $\text${h} minimum bias runs, after the correction to the PMT readout is introduced, so that PMT signals are discriminated at the leading edge in both the emulation and the hardware. The fiducialized $[0.2,1.9]$ µs beam gate window is highlighted.